An important part of customer service responsibilities of a telephone company relates to the speed with...

  1. An important part of customer service responsibilities of a telephone company relates to the speed with which troubles in residential service can be repaired. Suppose past data indicate that the likelihood is 0.8 that troubles in residential service can be repaired on the same day. For the first five troubles reported on a given day, what is the probability that less than 5 will be repaired on the same day?
a.

0.0012

b.

0.4096

c.

0.0003

d.

0.6723

e.

0.9988
Answer #1

Here, n = 5, p = 0.8, (1 - p) = 0.2 and x = 5
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X < 5).
P(X < 5) = (5C0 * 0.8^0 * 0.2^5) + (5C1 * 0.8^1 * 0.2^4) + (5C2 * 0.8^2 * 0.2^3) + (5C3 * 0.8^3 * 0.2^2) + (5C4 * 0.8^4 * 0.2^1)
P(X <5) = 0.0003 + 0.0064 + 0.0512 + 0.2048 + 0.4096
P(X < 5) = 0.6723